1.5Nos new camera installation
2.1Nos New PoE Switch Installation
3.1Nos New Onu Installation
4.CCTV Server Panel Installation
5.Optical Fiver Connectivity












1.5Nos new camera installation
2.1Nos New PoE Switch Installation
3.1Nos New Onu Installation
4.CCTV Server Panel Installation
5.Optical Fiver Connectivity












Checking and maintenance of 2 field pumps completed.
Due To Strom the optical fiber cable fall down near DTW-114 to 117 and chaira brize to DTW-218 now everything is normal
IRP-1 All sandfilter ultrasonic sensor cleaning Done. There was so much dust
Observation: Radio communication checking complete DTW 221.1. In DTW 221, We checked up to 10 minutes. RSSI Value (Dc volts) MAX: 2.700V, MIN:…
Problem: Shutdown time is ACB not open.Problem found: ACB Internal control unit not working found.Observation: DTW 203 UVR checked, When plug PTT UVR…